OpenStudy (anonymous):

Inverse of Ln(1+x) PS. Look at it carefully! ^^

OpenStudy (bahrom7893):

y = Ln(1+x) e^y = e^[Ln(1+x)] e^y = 1 + x

OpenStudy (bahrom7893):

x = e^y - 1

OpenStudy (anonymous):

Close, but no Bahroom...^^

OpenStudy (bahrom7893):

No? what's the answer, i think its just in a different format.

OpenStudy (anonymous):

It should be in term of y! xD

OpenStudy (bahrom7893):

wait what's in terms of y?

OpenStudy (bahrom7893):

it is in terms of y; x = f(y)

OpenStudy (anonymous):

It's y = Ln(x+1) and I want the inverse in term of y = .....

OpenStudy (bahrom7893):

No, the answer is: x = e^y + 1

OpenStudy (anonymous):

Nah...It's y = e^x -1

OpenStudy (bahrom7893):

oh okay well that is the inverse of my anwer lol, to make an inverse, simply change the spots of y and x

OpenStudy (anonymous):

Yup! xD and then you solve for y, cause the calculator doesn't do x! ^^ You're correct, but not complete....

OpenStudy (anonymous):

okay i think i get it

OpenStudy (anonymous):

notice the hesitance i have

OpenStudy (anonymous):

why are you looking at this hanh?